# Optimizing energy conversion with nonthermal resources in steady-state quantum devices

### v1

@@ III Optimizing energy conversion
To optimize the conversion of an available energy resource into a useful output, such as cooling power, we use the previously introduced performance quantifiers as optimization objectives. We have a general approach where the task is to achieve a generic output current, $I^{x}>0$. The goal is then to optimize different performance quantifiers, which are (A) the generic output currents themselves, (B) the efficiencies of their generation with generic input currents (Eq. (7)), (C) the precision of these output currents by the reduction of their noise (Eq. (8)), and finally (D) a combinations of these as a trade-off relation (Eq. (9)). The aim of this paper is to find the optimal transmission function of the two-terminal quantum-scattering setup for the performance quantifiers. This approach, based on scattering theory, provides insights to guide the design of energy-conversion experiments and sets bounds on achievable performance, whenever independent transmission channels are present in nanostructures [37, 4].
We begin with the straightforward optimization of the desired output current shown in Sec. III.1. As a next step, we optimize the other performance quantifiers under the constraint of a given output current. The strategies for finding the transmission functions for optimizing these different constrained performance quantifiers are analogous to each other and they are similar to variational principles using Lagrange multipliers [33, 11]. In Refs. [33, 11], the efficiency at given output power in a two-terminal device with thermal contacts was shown to be reached by having a boxcar-shaped transmission probability. Here, we extend this approach to systems with nonthermal distributions and optimize furthermore the efficiencies of generic input and output currents—including those where the current coefficient $x(\varepsilon)$ is nonlinear in energy such as in the entropy current—and of the precision, see also Ref. [34].
@@ III.2 Best efficiency at fixed output current
As a next step, we optimize the efficiency at a given output current. We therefore follow the strategy of Ref. [33, 11, 35], where it was shown that the transmission probability of a coherent, non-interacting two-terminal conductor with thermal contacts which maximizes the efficiency for a given power has a boxcar shape. However, we have already seen in Eq. (12) and in Fig. 2 for the optimization of the output current that the nonmonotonicity of the occupation probabilities and hence of the spectral currents leads to a more intricate shape of the optimal transmission probability.
@@
$I^{x}_{\gamma}(d_{\gamma})$ $=\delta\varepsilon d_{\gamma}x_{\gamma}(f_{\gamma}-g_{\gamma})$ (13a)
@@
Instead, we find a global minimum by considering the domain boundaries of $d_{\gamma}$. We therefore now ask, how the integrand behaves with variations of the transmission. If it is beneficial for the transmission function to increase for a certain energy, then the global minimum must be obtained when the transmission is equal to one for that point (and equal to zero in the opposite case). Explicitly, we set
$\left.\frac{\partial I^{y}}{\partial d_{\gamma}}\right|_{I^{x}_{\mathrm{fix}}}<0\longrightarrow d_{\gamma}=1$ (16a)
@@
We now take the continuous limit of the energy spectrum by letting $\delta\varepsilon\longrightarrow 0$ to find a condition for the full transmission function $\mathcal{D}(\varepsilon)$. The continuous equivalent of Eq. (14) is
@@
In analogy to the discrete case, we aim to extract the optimal transmission from minimizing $I^{y}$. Following usual calculus of variations [55], the current under variation of the transmission function and constraint $I^{x}_{\mathrm{fix}}$ reads,
@@
Here, variational calculus reaches an impasse; as previously argued there should not exist any local minimum with regards to $\mathcal{D}(\varepsilon)$, meaning that setting $\delta I^{y}|_{I^{x}_{\mathrm{fix}}}=0$ and solving like a regular variational problem yields nonsensical results. Instead, the continuous notion of a variation is combined with the knowledge from the discrete optimization by introducing an inequaltity: $\delta I^{y}|_{I^{x}_{\mathrm{fix}}}<0$. Furthermore, Eq. (15) implies that the energy-resolved inequality lets us find the optimal transmission, by saturating the bounds of $\mathcal{D}(\varepsilon)$ to either 0 or 1 depending on the sign. Similar to how one in caluculus of variations solves for the stationary point with the integrand, we find
@@
Figure 4: Finding the transmission function with crossing points. Purple lines are related to quantities of the nonthermal reservoir, blue lines to quantities of the cold one. (a) Two distribution functions, where red indicates regions where the second factor in the argument of Eq. (12) or equivalently of Eq. (21) is positive. (b) The two current coefficients entering the first factor of Eq. (21); $y(\varepsilon)$ gives the entropy current into the nonthermal bath and $\lambda x(\varepsilon)$ the entropy current out of the cold bath, weighted by the Lagrange multiplier. Panel (c) illustrates where in the energy spectrum the two factors are negative. In the fainter red areas, the step-function in Eq. (21) yields zero. In the white and in the darker red area, the factors in the argument of Eq. (21) are both positive or both negative, such that the transmission function is set to 1.
We find an insightful interpretation of $\lambda$ in terms of a characteristic efficiency when inspecting the condition (III.2) further. It is readily rewritten as
@@ III.3 Lowest noise at fixed output current
Technically, there is an important difference between the minimization of the noise and the optimization of average currents: the noise has a quadratic contribution in $\mathcal{D}(\varepsilon)$. Note however that the shot-noise term $S^{x}_{\mathrm{shot}}$, which contains all quadratic dependence in $\mathcal{D}(\varepsilon)$, is never negative (it hence always increases the noise) and that the factor $\mathcal{D}(\varepsilon)(1-\mathcal{D}(\varepsilon))$ implies that this term vanishes whenever $\mathcal{D}(\varepsilon)=0,1$. The thermal-like noise in turn is linear in the transmission probability and we can hence apply the same procedure to minimize it as applied in the previous section for the optimization of the efficiency. Since the shot noise is suppressed for a boxcar-shaped transmission, and since the thermal noise is linear in $\mathcal{D}(\varepsilon)$, we can anticipate that the optimal transmission for the minimization of noise will be a series of boxcars. Thus, the optimization can be done just on $S^{x}_{\mathrm{thermal}}$ in a similar fashion as for the efficiencies above. More rigorous derivations of the fact that it is sufficient to minimize the thermal noise can be found in Appendix B and in Ref. [34]. Briefly, the terms quadratic in $\mathcal{D}(\varepsilon)$ are concave, such that the minimum points of its derivative are at the boundaries like in the linear case, see Fig. 3(b).
@@
which is hence the function that minimizes the noise of a generic fixed output current. The Lagrange multiplier $\lambda$ is again found from the constraint $I^{x}\equiv I^{x}_{\mathrm{fix}}$. In the case where the contacts are thermal and the noise of charge or energy currents is considered, the result of Eq. (31) reduces to the recent result on precision optimization of Ref. [34].
@@
$\mathcal{D}^{S}_{\mathrm{opt}}(\varepsilon)=\Theta(\mathrm{NSR}^{\mathrm{char}}_{\mathrm{max}}-\mathrm{NSR}^{\mathrm{char}}(\varepsilon))$ (33)
@@
This ratio is bounded from below by 0, but can be arbitrarily large. Since $i^{x}(\varepsilon)>0$ is considered beneficial, only $\mathrm{NSR}^{\mathrm{char}}_{\mathrm{max}}>0$ gives a positive output current. But $i^{x}(\varepsilon)$ can be arbitrarily close to zero while positive, which always happens near crossing points between the distributions, see Eq. (4). Thus, $\mathrm{NSR}^{\mathrm{char}}_{\mathrm{max}}\to\infty$ as the $\mathcal{D}^{S}_{\mathrm{opt}}(\varepsilon)$ approaches the transmission function for maximum output, Eq (12). Furthermore, $\mathcal{D}^{S}_{\mathrm{opt}}(\varepsilon)$ tends to exclude areas close to where $i^{x}(\varepsilon)\to 0$.
@@ III.4 Best precision-efficiency trade-off for fixed output current
We have shown in the previous sections how to identify transmission probabilities that result in the best efficiency and how to identify transmission probabilities that result in the lowest noise at a given output current. Our previous arguments show that also for nonthermal distributions, the transmission probabilities that optimize these two goals need to be different, see, e.g. Fig. 7(f) compared to (g). Indeed, while the efficiency is best when regions of finite transmission select the crossing points between distribution functions [9, 10], the precision is best when the distribution functions differ maximally in the transport window selected by the transmission probability [56]. As a next step, we therefore aim to optimize the trade-off between the two. Such a trade-off can for example be formulated in terms of a thermodynamic uncertainty relation (TUR), see Eq. (9).
@@
where the differentiation with regards to $d_{\gamma}$ picks out one energy slice, while the noise $S^{x}$ and the current $I^{y}$, which occur as prefactors of the derivatives in Eq. (34), are sums over all slices. Then the procedure follows as in Sec. III.2, with $d_{\gamma}$ being set to either 0 or 1 depending on the sign of the derivatives, and with $S^{x}$ replaced by $S^{x}_{\mathrm{thermal}}$ as in Sec. III.3. The key difference is that the constants $S^{x}$ and $I^{y}$ are not arbitrary but must be solved self-consistently, similarly to how $\lambda$ must be solved for a specific fixed output current $I^{x}$.
@@
$\mathcal{D}(\varepsilon)=\Theta\Big{(}-2\alpha I^{y}s^{x}(\varepsilon))-2(1-\alpha)S^{x}i^{y}(\varepsilon)+\lambda i^{x}(\varepsilon)\Big{)}\ .$ (39)
@@
III.5 Optimization with regards to a second constant
We have until here optimized the performance of the energy-conversion process by tuning the transmission function of the conductor. In many cases, it might be relevant to optimize with regards to another parameter in addition to the transmission function. For example, for thermal reservoirs it is useful to consider the potential difference as a variable, as it can typically be externally modified. In earlier results [33, 11] for the optimization of efficiency at fixed power, the variation in terms of the potential difference was directly used when determining that the optimal transmission function was a boxcar. In the following we present a method that mirrors very closely the procedure in Ref. [11], but considering any arbitrary second variable.
We hence choose a second variable $\xi$, which is independent of the energy $\varepsilon$. We consider as our goal to maximize the efficiency with a fixed output current $I^{x}_{\mathrm{fix}}$, like in Sec. III.2, but now by varying both $\mathcal{D}(\varepsilon)$ and $\xi$. The approach is to solve for the unknown $\lambda$ by introducing a second constraint equation. Since $\xi$ is not energy-dependent, the variations of the currents with regards to it are direct partial derivatives, and as such they are not energy-resolved. We find
$\delta_{\mathcal{D}}I^{y}|_{\xi}-\lambda\delta_{D}I^{x}|_{\xi}$ $<$ $0$ (40a)
$\left.\frac{\partial I^{y}}{\partial\xi}\right|_{\mathcal{D}}-\left.\lambda\frac{\partial I^{x}}{\partial\xi}\right|_{\mathcal{D}}$ $=$ $0$ (40b)
@@ III.5 Optimization with regards to a second constant
For terminals with thermal distributions and with $\xi\equiv\Delta\mu$, this condition coincides with the result of Ref. [11]. Effectively, in the situation where the efficiency is optimized by tuning $\mu$ in addition to $\mathcal{D}(\varepsilon)$, the unknown constant $\lambda$ in previous sections is replaced by the derivatives of the currents with regards to the second variable, evaluated at fixed transmission. In other words, the optimal $\xi$ for a certain value of $I^{x}$ is found at the point where the Lagrange multiplier $\lambda$ for a given $\xi$ and the fraction of derivatives $\frac{\partial I^{y}/\partial\xi|_{\mathcal{D}}}{\partial I^{x}/\partial\xi|_{\mathcal{D}}}$ coincide. We visualize this in Fig. 6 for the example of finding the optimal efficiency for a certain fixed current (here $I^{x}\to I^{\Sigma,\mathrm{L}}$) over a range of $\mu$, which here represents a shift in the electrochemical potential of the cold reservoir compared to the zero of energy with respect to which the nonthermal distribution in terminal R is defined. Panel (a) shows the result of the efficiency optimized “brute-force” with respect to $\mu$, by finding the optimal transmission function for the same $I^{\Sigma,\mathrm{L}}_{\mathrm{fix}}$ over a range of $\mu$. This result perfectly coincides with the crossing of $\lambda(\mu)$ and the calculated fraction of derivatives, as sketched in Fig. 6(b).
However, note that the constraint in Eq. (40b) is an equality, as opposed to an inequality like Eq. (40a). This means that there can be several solutions that fulfill the criterion $I^{x}_{\mathrm{fix}}$, if $I^{y}$ has several extrema with regards to $\xi$ and optimized transmission functions. This is illustrated in panels (c) and (d) in Fig. 6. In panel (c), there is a local maximum and minimum, in which Eq. (40b) is fulfilled. This is clear in panel (d), where $\lambda$ is shown to coincide with the fraction of derivatives at both points. Thus, the more favorable solution should be selected, if co-existing solutions exist.
@@
Figure 6: Optimized efficiencies for the fixed left entropy current $I^{\Sigma,\mathrm{L}}=0.15\>k_{\mathrm{B}}^{2}T/h$ by varying the right entropy current $I^{\Sigma,\mathrm{R}}$ over a range of $\mu\in[0,10]\ k_{\mathrm{B}}T$, between the cold thermal distribution in L and a nonthermal distribution with a dip-peak structure. (a) shows the calculated efficiencies with optimal transmission functions for each $\mu$, with the maximal efficiency found in $\mu=5.43\ k_{\mathrm{B}}T$. In (b), the light purple curve shows $1/\lambda$ associated with the optimal transmission functions at each $\mu$ and the dark purple, dashed curve is the calculated fraction of partial derivatives as found in Eq. (41). (c) and (d) are zooms of the upper panels, picking our the region $\mu\in[1,2]\ k_{\mathrm{B}}T$ where additional extrema occur.
@@
$\lambda=\frac{2\alpha I^{y}\partial S^{x}/\partial\xi+2(1-\alpha)S^{x}\partial I^{y}/\partial\xi}{\partial I^{x}/\partial\xi}\ .$ (42)
This is hence an analogous problem to the one we solved in Sec. III.4, but now $\lambda$ is determined by the derivatives of the currents and fluctuations with regards to the second parameter $\xi$, at a fixed transmission function. For the specific values of $\alpha=0,1/2,1$, corresponding to an optimization of efficiency, TUR-like trade-off, and low noise-to-signal ratio at fixed output current, one finds
$\alpha$ $=$ $0\rightarrow\lambda=\frac{\partial I^{y}/\partial\xi}{\partial I^{x}/\partial\xi}$ (43a)
$\alpha$ $=$ $1/2\rightarrow\lambda=\frac{\partial S^{x}/\partial\xi}{\partial I^{x}/\partial\xi}$ (43b)
$\alpha$ $=$ $1\rightarrow\lambda=\frac{\partial(S^{x}I^{y})/\partial\xi}{\partial I^{x}/\partial\xi}\ .$ (43c)

### v2

@@ III Optimizing energy conversion
To optimize the conversion of an available energy resource into a useful output, such as cooling power, we use the previously introduced performance quantifiers as optimization objectives. We have a general approach where the task is to achieve a generic output current, $I^{x}>0$. The goal is then to optimize different performance quantifiers, which are (A) the generic output currents themselves, (B) the efficiencies of their generation with generic input currents (Eq. (8)), (C) the precision of these output currents by the reduction of their noise (Eq. (9)), and finally (D) a combinations of these as a trade-off relation (Eq. (11)). The aim of this paper is to find the optimal transmission function of the two-terminal quantum-scattering setup for the performance quantifiers. This approach, based on scattering theory, provides insights to guide the design of energy-conversion experiments and sets bounds on achievable performance, whenever independent transmission channels are present in nanostructures [VanHouten1992, Benenti2017Jun].
We begin with the straightforward optimization of the desired output current shown in Sec. III.1. As a next step, we optimize the other performance quantifiers under the constraint of a given output current. The strategies for finding the transmission functions for optimizing these different constrained performance quantifiers are analogous to each other and they are similar to variational principles using Lagrange multipliers [Whitney2014Apr, Whitney2015]. In Refs. [Whitney2014Apr, Whitney2015], the efficiency at given output power in a two-terminal device with thermal contacts was shown to be reached by having a boxcar-shaped transmission probability. Here, we extend this approach to systems with nonthermal distributions and optimize furthermore the efficiencies of generic input and output currents—including those where the current coefficient $x(\varepsilon)$ is nonlinear in energy such as in the entropy current—and of the precision, see also Ref. [Landi2025].
@@ III.2 Best efficiency at fixed output current
As a next step, we optimize the efficiency at a given output current. We therefore follow the strategy of Ref. [Whitney2014Apr, Whitney2015, Chrirou2025Jul], where it was shown that the transmission probability of a coherent, non-interacting two-terminal conductor with thermal contacts which maximizes the efficiency for a given power has a boxcar shape. However, we have already seen in Eq. (13) and in Fig. 2 for the optimization of the output current that the nonmonotonicity of the occupation probabilities and hence of the spectral currents leads to a more intricate shape of the optimal transmission probability.
@@
$I^{x}_{\gamma}(d_{\gamma})$ $=\delta\varepsilon d_{\gamma}x_{\gamma}(f_{\gamma}-g_{\gamma}),$ (14a)
@@
Instead, we find a global minimum by considering the domain boundaries of $d_{\gamma}$. We therefore now ask how the integrand behaves with variations of the transmission. If it is beneficial for the transmission function to increase for a certain energy, then the global minimum must be obtained when the transmission is equal to one for that point (and equal to zero in the opposite case). Explicitly, we set
$\left.\frac{\partial I^{y}}{\partial d_{\gamma}}\right|_{I^{x}_{\mathrm{fix}}}<0\longrightarrow d_{\gamma}=1,$ (17a)
@@
We now take the continuous limit of the energy spectrum by letting $\delta\varepsilon\to 0$ to find a condition for the entire transmission function $\mathcal{D}(\varepsilon)$. The continuous equivalent of Eq. (15) is
@@
In analogy to the discrete case, we aim to extract the optimal transmission from minimizing $I^{y}$. Following usual calculus of variations [Arfken], the current under variation of the transmission function and constraint $I^{x}_{\mathrm{fix}}$ reads,
@@
Here, variational calculus reaches an impasse; as previously argued there should not exist any local minimum with respect to $\mathcal{D}(\varepsilon)$, meaning that setting $\delta I^{y}|_{I^{x}_{\mathrm{fix}}}=0$ and solving like a regular variational problem yields nonsensical results. Instead, the continuous notion of a variation is combined with the knowledge from the discrete optimization by introducing an inequaltity: $\delta I^{y}|_{I^{x}_{\mathrm{fix}}}<0$. Furthermore, Eq. (16) implies that the energy-resolved inequality lets us find the optimal transmission, by saturating the bounds of $\mathcal{D}(\varepsilon)$ to either 0 or 1 depending on the sign. Similar to how one in caluculus of variations solves for the stationary point with the integrand, we find
@@
Figure 4: Finding the transmission function with crossing points. Purple lines are related to quantities of the nonthermal contact, blue lines to quantities of the cold one. (a) Two distribution functions, where red indicates regions where the second factor in the argument of Eq. (22) is positive. (b) The two current coefficients entering the first factor of Eq. (22); $y(\varepsilon)$ gives the entropy current into the nonthermal bath and $\lambda x(\varepsilon)$ the entropy current out of the cold bath, weighted by the Lagrange multiplier. Panel (c) illustrates where in the energy spectrum the two factors are negative. In the fainter red areas, the step-function in Eq. (22) yields zero. In the white and in the darker red area, the factors in the argument of Eq. (22) are both positive or both negative, such that the transmission function is set to 1.
We find an insightful interpretation of $\lambda$ in terms of a characteristic efficiency when inspecting the condition in Eq. (III.2) further. It is readily rewritten as
@@ III.3 Lowest noise at fixed output current
Technically, there is an important difference between the minimization of the noise and the optimization of average currents: the noise has a quadratic contribution in $\mathcal{D}(\varepsilon)$. Note however that the shot-noise term $S^{x}_{\mathrm{shot}}$, which contains all quadratic dependence in $\mathcal{D}(\varepsilon)$, is never negative (it hence always increases the noise) and that the factor $\mathcal{D}(\varepsilon)(1-\mathcal{D}(\varepsilon))$ implies that this term vanishes whenever $\mathcal{D}(\varepsilon)=0,1$. The thermal-like noise in turn is linear in the transmission probability and we can hence apply the same procedure to minimize it as applied in the previous section for the optimization of the efficiency. Since the shot noise is suppressed for a boxcar-shaped transmission, and since the thermal noise is linear in $\mathcal{D}(\varepsilon)$, we can anticipate that the optimal transmission for the minimization of noise will be a series of boxcars. Thus, the optimization can be done just on $S^{x}_{\mathrm{thermal}}$ in a similar fashion as for the efficiencies above. More rigorous derivations of the fact that it is sufficient to minimize the thermal noise can be found in Appendix B and in Ref. [Landi2025]. Briefly, the terms quadratic in $\mathcal{D}(\varepsilon)$ are concave, such that the minimum points of its derivative are at the boundaries like in the linear case, see Fig. 3(b).
@@
which is hence the function that minimizes the noise of a generic fixed output current. The Lagrange multiplier $\lambda$ is again found from the constraint $I^{x}\equiv I^{x}_{\mathrm{fix}}$. In the case where the contacts are thermal and the noise of charge or energy currents is considered, the result of Eq. (32) reduces to the recent result on precision optimization of Ref. [Landi2025].
@@
$\mathcal{D}^{S}_{\mathrm{opt}}(\varepsilon)=\Theta(\mathrm{NSR}^{\mathrm{char}}_{\mathrm{max}}-\mathrm{NSR}^{\mathrm{char}}(\varepsilon)).$ (34)
@@
This ratio is bounded from below by 0, but can be arbitrarily large. Since $i^{x}(\varepsilon)>0$ is considered beneficial, only $\mathrm{NSR}^{\mathrm{char}}_{\mathrm{max}}>0$ gives a positive output current. But $i^{x}(\varepsilon)$ can be arbitrarily close to zero while positive, which always happens near crossing points between the distributions, see Eq. (4). Thus, $\mathrm{NSR}^{\mathrm{char}}_{\mathrm{max}}\to\infty$ as the $\mathcal{D}^{S}_{\mathrm{opt}}(\varepsilon)$ approaches the transmission function for maximum output, Eq. (13). Furthermore, $\mathcal{D}^{S}_{\mathrm{opt}}(\varepsilon)$ tends to exclude areas close to where $i^{x}(\varepsilon)\to 0$.
@@ III.4 Best precision-efficiency trade-off for fixed output current
We have shown in the previous sections how to identify transmission probabilities that result in the best efficiency and how to identify transmission probabilities that result in the lowest noise at a given output current. Our previous arguments show that also for nonthermal distributions, the transmission probabilities that optimize these two goals need to be different, see in addition the discussion in IV. Indeed, while the efficiency is best when regions of finite transmission select the crossing points between distribution functions [Hicks1993May, Mahan1996Jul], the precision is best when the distribution functions differ maximally in the transport window selected by the transmission probability [Palmqvist2024]. As a next step, we therefore aim to optimize the trade-off between the two. Such a trade-off can for example be formulated in terms of a thermodynamic uncertainty relation (TUR), see Eq. (11).
@@
where the differentiation with respect to $d_{\gamma}$ picks out one energy slice, while the noise $S^{x}$ and the current $I^{y}$, which occur as prefactors of the derivatives in Eq. (35), are sums over all slices. Then the procedure follows as in Sec. III.2, with $d_{\gamma}$ being set to either 0 or 1 depending on the sign of the derivatives, and with $S^{x}$ replaced by $S^{x}_{\mathrm{thermal}}$ as in Sec. III.3. The key difference is that the constants $S^{x}$ and $I^{y}$ are not arbitrary but must be solved self-consistently, similarly to how $\lambda$ must be solved for a specific fixed output current $I^{x}$.
@@
$\mathcal{D}_{\mathrm{opt}}^{\alpha}(\varepsilon)=\Theta(-2\alpha I^{y}s^{x}(\varepsilon)-2(1-\alpha)S^{x}i^{y}(\varepsilon)+\lambda i^{x}(\varepsilon)).$ (40)
@@
III.5 Optimization with respect to a second constant
We have until here optimized the performance of the energy-conversion process by tuning the transmission function of the conductor. In many cases, it might be relevant to optimize with respect to another parameter in addition to the transmission function. For example, for thermal reservoirs it is useful to consider the potential difference as a variable, as it can typically be externally modified. In earlier results [Whitney2014Apr, Whitney2015] for the optimization of efficiency at fixed power, the variation in terms of the potential difference was directly used when determining that the optimal transmission function was a boxcar. In the following we present a method that mirrors very closely the procedure in Ref. [Whitney2015], but considering any arbitrary second variable.
We hence choose a second variable $\xi$, which is independent of the energy $\varepsilon$. We consider as our goal the maximization of the efficiency with a fixed output current $I^{x}_{\mathrm{fix}}$, like in Sec. III.2, but now by varying both $\mathcal{D}(\varepsilon)$ and $\xi$. The approach is to solve for the unknown $\lambda$ by introducing a second constraint equation. Since $\xi$ is not energy-dependent, the variations of the currents with respect to it are direct partial derivatives, and as such they are not energy-resolved. We find
$\delta_{\mathcal{D}}I^{y}|_{\xi}-\lambda\delta_{D}I^{x}|_{\xi}$ $<$ $0,$ (41a)
$\left.\frac{\partial I^{y}}{\partial\xi}\right|_{\mathcal{D}}-\left.\lambda\frac{\partial I^{x}}{\partial\xi}\right|_{\mathcal{D}}$ $=$ $0.$ (41b)
@@ III.5 Optimization with respect to a second constant
For terminals with thermal distributions and with $\xi\equiv\Delta\mu$, this condition coincides with the result of Ref. [Whitney2015]. Effectively, in the situation where the efficiency is optimized by tuning $\mu$ in addition to $\mathcal{D}(\varepsilon)$, the unknown constant $\lambda$ in previous sections is replaced by the derivatives of the currents with respect to the second variable, evaluated at fixed transmission function defined by the full transmission condition. In other words, the optimal $\xi$ for a certain value of $I^{x}$ is found at the point where the Lagrange multiplier $\lambda$ for a given $\xi$ and the fraction of derivatives $\frac{\partial I^{y}/\partial\xi|_{\mathcal{D}}}{\partial I^{x}/\partial\xi|_{\mathcal{D}}}$ coincide. We visualize this in Fig. 6 for the example of finding the optimal efficiency for a certain fixed current (here $I^{x}\to I^{\Sigma,\mathrm{L}}$) over a range of $\mu$, which here represents a shift in the electrochemical potential of the cold reservoir compared to the zero of energy with respect to which the nonthermal distribution in terminal R is defined. Panel (a) shows the result of the efficiency optimized by “brute-force” with respect to $\mu$, by finding the optimal transmission function for the same $I^{\Sigma,\mathrm{L}}_{\mathrm{fix}}$ over a range of $\mu$. This result perfectly coincides with the crossing of $\lambda(\mu)$ and the calculated fraction of derivatives, as sketched in Fig. 6(b).
However, note that the constraint in Eq. (41b) is an equality, as opposed to an inequality like Eq. (41a). This means that there can be several solutions that fulfill the criterion $I^{x}_{\mathrm{fix}}$, if $I^{y}$ has several extrema with respect to $\xi$ and optimized transmission functions. This is illustrated in panels (c) and (d) in Fig. 6. In panel (c), there is a local maximum and minimum, in which Eq. (41b) is fulfilled. This is clear in panel (d), where $\lambda$ is shown to coincide with the fraction of derivatives at both points. Thus, the more favorable solution should be selected, if co-existing solutions exist.
@@
Figure 6: Optimized efficiencies for the fixed left entropy current $I^{\Sigma,\mathrm{L}}=0.15\>k_{\mathrm{B}}^{2}T/h$ by varying the right entropy current $I^{\Sigma,\mathrm{R}}$ over a range of $\mu\in[0,10]\ k_{\mathrm{B}}T$, between the cold thermal distribution in L and a nonthermal distribution with a dip-peak structure. (a) shows the calculated efficiencies with optimal transmission functions for each $\mu$, with the maximal efficiency found in $\mu=5.43\ k_{\mathrm{B}}T$. In (b), the light purple curve shows $1/\lambda$ associated with the optimal transmission functions at each $\mu$ and the dark purple, dashed curve is the calculated fraction of partial derivatives as found in Eq. (42). (c) and (d) are zooms of the upper panels, picking our the region $\mu\in[1,2]\ k_{\mathrm{B}}T$ where additional extrema occur (indicated by black veritcal lines).
@@
$\lambda=\frac{2\alpha I^{y}\partial S^{x}/\partial\xi|_{\mathcal{D}}+2(1-\alpha)S^{x}\partial I^{y}/\partial\xi|_{\mathcal{D}}}{\partial I^{x}/\partial\xi|_{\mathcal{D}}}\ .$ (43)
This is hence an analogous problem to the one we solved in Sec. III.4, but now $\lambda$ is determined by the derivatives of the currents and fluctuations with respect to the second parameter $\xi$, at a fixed transmission function. For the specific values of $\alpha=0,1/2,1$, corresponding to an optimization of efficiency, TUR-like trade-off, and low noise-to-signal ratio at fixed output current, one finds
$\alpha$ $=$ $0\rightarrow\lambda=\frac{\partial I^{y}/\partial\xi|_{\mathcal{D}}}{\partial I^{x}/\partial\xi|_{\mathcal{D}}},$ (44a)
$\alpha$ $=$ $1/2\rightarrow\lambda=\frac{\partial S^{x}/\partial\xi|_{\mathcal{D}}}{\partial I^{x}/\partial\xi|_{\mathcal{D}}},$ (44b)
$\alpha$ $=$ $1\rightarrow\lambda=\frac{\partial(S^{x}I^{y})/\partial\xi|_{\mathcal{D}}}{\partial I^{x}/\partial\xi|_{\mathcal{D}}}.$ (44c)